### v2

# A Peek into Token Bias: Large Language Models Are Not Yet Genuine Reasoners

Abstract
This study introduces a hypothesis-testing framework to assess whether large language models (LLMs) possess genuine reasoning abilities or primarily depend on token bias. We go beyond evaluating LLMs on accuracy; rather, we aim to investigate their token bias in solving logical reasoning tasks. Specifically, we develop carefully controlled synthetic datasets, featuring conjunction fallacy and syllogistic problems. Our framework outlines a list of hypotheses where token biases are readily identifiable, with all null hypotheses assuming genuine reasoning capabilities of LLMs. The findings in this study suggest, with statistical guarantee, that most LLMs still struggle with logical reasoning. While they may perform well on classic problems, their success largely depends on recognizing superficial patterns with strong token bias, thereby raising concerns about their actual reasoning and generalization abilities. Codes and data are open-sourced at https://github.com/bowen-upenn/llm_token_bias.
A Peek into Token Bias: Large Language Models Are Not Yet
Genuine Reasoners
Bowen Jiang1, 2, Yangxinyu Xie1, 2, Zhuoqun Hao1, Xiaomeng Wang1, Tanwi Mallick2, Weijie J. Su1, Camillo J. Taylor1, Dan Roth1    University of Pennsylvania1   Argonne National Laboratory2 Philadelphia, PA, 19104, USA   Lemont, IL, 60439, USA {bwjiang@seas, xinyux@wharton, zhuoqunh@sas, xwang1@wharton}.upenn.edu, tmallick@anl.gov, {suw@wharton, cjtaylor@seas, danroth@seas}.upenn.edu
1 Introduction
Large language models (LLMs) have achieved remarkable progress in understanding and generating human-like text, triggering growing interest in the LLMs’ theory of minds Kosinski (2023); Jamali et al. (2023); Bubeck et al. (2023) and decision-making abilities Lyu et al. (2023); Prasad et al. (2023); Jiang et al. (2024a, b); Xie et al. (2024). However, there is ongoing debate about whether LLMs possess genuine reasoning capabilities, as evidence suggests that the performance of LLMs on reasoning tasks is correlated with how much the input’s semantic content supports a correct logical inference Dasgupta et al. (2022); Li et al. (2023). Should valid reasoning be applied, such a correlation would not exist, since a genuine reasoner should be able to derive the correct inference regardless of the context.
In this paper, we formalize this observation and say that an LLM is subject to token bias in a reasoning task if systematic changes to some or all tokens in the task descriptions - while keeping the underlying logic intact - allow us to predict the direction of the shift in the model’s output. A strong token bias suggests that the model is relying on superficial patterns in the input rather than truly understanding the underlying reasoning task. This could lead to brittle performance that fails to generalize well to novel examples and phrasings encountered in the wild, which could differ from the spurious patterns the model may have overfitted to during training.
Figure 1: We illustrate token bias using the classic "twenty-five horses" problem in graph theory. The top two sub-figures, generated by GPT-4o for illustration purposes only [^2] Interestingly, when we prompted GPT-4o to generate an image of "lop-eared bunnies", the model exhibited a visual token bias by depicting bunnies with four ears — both lop and erect — implying it associated the term "bunnies" with the presence of erect ears in images, without a genuine and logical understanding of a bunny’s physical reality. We move the exploration of visual token bias to the future work., demonstrate the concept by altering the name "horses" to "bunnies", irrelevant to the problem’s underlying logic. The bottom two sub-figures show experimental results in GPT-4 and Claude, where performance significantly drops due to perturbations in animal names and numbers. In these plots, "Original" refers to the unaltered "twenty-five horses" problem, "random_animals" alters only the animal names, and "random" alters both names and numbers. We observe ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}$ with statistical significance, meaning that there are more instances where the original problem is solved correctly while the perturbed problem is solved incorrectly, compared to the reverse. As a result, our hypothesis testing confirms token bias in this scenario.
Figure 2: An illustration of the overall framework. We generate synthetic data, perform systematic token perturbations, and evaluate an LLM for comparative studies. The resulting contingency table, where A-D are integer values of counts, allows for subsequent statistical tests.
Figure 3: What is token bias? Here is another example exhibited by GPT-4. On the left, GPT-4 correctly identifies the conjunction fallacy and answers the question correctly, given the classical Linda Problem as the one-shot exemplar. On the right, however, the exemplar is rephrased by altering "Linda" to "Bob" while keeping the same logic, which surprisingly confuses the model.
We explore several well-known logical fallacy problems from the cognitive science literature Tversky and Kahneman (1983); Kahneman (2011), which provide a clear playground for assessing the reasoning capabilities of LLMs. Figure 2 and 3 depict two kinds of token biases found in our testing framework, where the model may be overfitting to specific tokens commonly found in classic problem statements. Since we observe many cases where state-of-the-art LLMs like GPT-4 Achiam et al. (2023) successfully identify logical fallacies under certain settings, we highlight the urgent need for a framework to tease out whether LLMs apply genuine reasoning or merely exploit token bias for their improved performance.
This work reconceptualizes the evaluation of reasoning capabilities into a general and rigorous statistical testing framework. As shown in Figure 2, it comprises three critical components: synthetic data generation, token perturbation, and statistical hypothesis testing. This general framework is designed to bypass the complications of evaluation set contamination Zhou et al. (2023); Ravaut et al. (2024), leverage insights and tools from controlled experiments, and draw statistically valid conclusions.
Our study is unique from existing work Gou et al. (2023); Suri et al. (2024); Mukherjee and Chang (2024); Wang et al. (2024) in two folds. First, we are not evaluating the overall accuracy of LLMs in identifying different logical fallacies. Instead, our focus is on token bias. Although there are always more types of logical fallacies, we take the conjunction fallacy, syllogistic fallacy, and the "twenty-five horses" problem in graph theory as quintessential examples, which exhibit strong token biases that are more readily identifiable in their problem statements. By identifying and perturbing these specific tokens, we can induce predictable shifts in LLM responses. Second, we recognize that cognitive biases often emerge in implicit forms in real-life scenarios, so relying on engineering fine-grained prompts  Gou et al. (2023); Yao et al. (2024); Besta et al. (2024) to make LLMs identify specific logical fallacies is impractical for general-purpose user applications. As a result, we only leverage common prompting techniques that are sufficient to provide robust statistical evidence.
Comprehensive experiments on both commercial and open-sourced LLMs on large-scale synthetic datasets uncover a critical insight: it is the token bias that contributes the most to performance improvements in reasoning tasks, if any, rather than genuine advances in reasoning capabilities.
2 The General Framework
Our framework is summarized in Figure 2. This general framework is grounded on the premise that for a given reasoning task, a capable reasoning agent will consistently reach the same conclusion regardless of how the task is framed, as long as the underlying logic remains the same (Hastie and Dawes, 2009). This assumption lays the foundation of our null hypothesis, $\boldsymbol{H_{0}}$. In our setup, if an agent consistently applies reasoning in its decision-making process, the only source of failure should be the procedural mistakes during the agent’s abstract reasoning steps, which we assume to come up in an i.i.d. fashion. Our general framework contains three major parts as follows.
Synthetic Data Generation
Once the underlying logic of a reasoning task is defined, we create an algorithm to generate a synthetic dataset with $n$ samples. While it is helpful to leverage LLMs for linguistic coherence in the process, the data generation should be carefully controlled, utilizing information from real-world data or established datasets to mitigate potential biases from purely AI-generated texts. The process begins with the creation of a curated list of entities, encompassing diverse names, genders, ages, occupations, cultural backgrounds, and events where applicable, along with a textual template that dictates the structure of the task description. By sampling from this list, we generate task descriptions that maintain the integrity and novelty of the dataset. This method ensures that while the LLM of interest might be familiar with the individual entities, it has never seen the specific combinations of these entities and narratives, thus bypassing data contamination.
The following example illustrates one approach we leverage to generate synthetic conjunction fallacy questions. We randomly sample a commonsense story curated by Mostafazadeh et al. (2016) and convert it into the following prompt: Your task is to complete the last sentence of the following problem to create a conjunction fallacy quiz:
Michelle was extremely hungry. She opened the refrigerator to find nothing. Which is more likely?
(a) Michelle would likely buy food at the grocery store.
(b) Michelle would likely buy food at the grocery store because
We expect the LLM to complete the story by providing us with a plausible reason after "because", such as "she found nothing to eat at home". Irrespective of the LLM’s completion, option (b) contains a conjunction of two events so it should always be viewed less likely.
The synthetic dataset can be dynamically generated on the fly, precluding its prior existence in any training datasets. It also allows the algorithm designers to control the dataset size, efficiently scaling their data based on the sample size required to acheive statistical validity.
Token Perturbation
We posit that if the LLM primarily relies on token bias, its performance on reasoning tasks will consistently improve (or degrade) as we alter some tokens in a systematic manner. This process of token perturbation generates $n$ matched pairs of samples, enabling us to evaluate the LLM on both the original and perturbed datasets and create a $2\times 2$ contingency table below, where $n=n_{11}+n_{12}+n_{21}+n_{22}$.
Perturbed
Correct Wrong
Original Correct $n_{11}$ $n_{12}$
Wrong $n_{21}$ $n_{22}$
Table 1: A template for the contingency table. We follow the notations in this table to define $\pi_{12}$ and $\pi_{21}$ in the next paragraph for hypothesis testing.
Statistical Hypothesis Testing for Matched Pairs
In our context, we wish to decide whether or not some hypothesis concerning whether an agent reasons consistently is correct. The choice here lies between two decisions: accepting or rejecting the hypothesis. The decision procedure is called hypothesis testing (Lehmann et al., 1986). Throughout our discussion, we use $\boldsymbol{H_{0}}$ to denote the null hypothesis and $\boldsymbol{H_{a}}$ the alternative hypothesis.
For each of the $n$ matched pairs, let $\pi_{ab}$ denote the underlying probability of outcome $a$ for the original sample and outcome $b$ for the perturbed sample. In other words, for any nonnegative integer $m\leq n$,
$\mathbb{P}(n_{ab}=m)=\binom{n}{m}\pi_{ab}^{m}(1-\pi_{ab})^{n-m}$ (1)
As $n_{ab}$ counts the number of such pairs, $n_{ab}/n$ is the sample proportion, which is a consistent estimate of $\pi_{ab}$. The null hypothesis assumes the marginal homogeneity for binary matched pairs, i.e. $\pi_{12}=\pi_{21}$. For small samples, we can apply an exact test conditioned on $n^{*}=n_{21}+n_{12}$ Mosteller (1952); Agresti (2012). Under $\boldsymbol{H_{0}},n_{21}$ follows a $\operatorname{binomial}(n^{*},1/2)$ distribution, and the corresponding $p$-value is the binomial tail probability. As a rule of thumb, when $n^{*}>10$, the reference binomial distribution is approximately normal, and we can compute the standardized normal test statistics $z_{0}=(n_{21}-n_{12})/\sqrt{n_{21}+n_{12}}$, which is identical to the McNemar statistic McNemar (1947). To test the same hypotheses for a group of models, we apply the Benjamini-Hochberg Procedure Benjamini and Hochberg (1995) to control the false discovery rate at a predetermined significance level $\alpha$.
3 A Peek into Token Bias
This section outlines the detailed hypotheses in our statistically inspired framework. We aim to determine whether LLMs are capable of genuine reasoning or whether they rely heavily on token biases. According to the principle of invariance in rational decision-making Tversky and Kahneman (1981, 1988), the preferences of a rational reasoning agent should remain unaffected by the framing of equivalent decision problems.
In a broader interpretation of invariance, we assess whether alterations in seemingly irrelevant tokens, such as name entities in problem narratives that are unrelated to the underlying logic, influence the outcomes of reasoning. A true reasoner should effectively navigate through reasoning tasks without being influenced by trivial changes in content that do not impact the fundamental logical structure. We propose a series of hypotheses, where the null hypothesis assumes the presence of a genuine reasoner. For each hypothesis, we identify specific tokens that may carry strong biases under their problem settings, and systematically alter these tokens to test their impact, while maintaining the integrity of the underlying logical structure.
3.1 Preliminaries
In this work, we integrate the conjunction fallacy and syllogistic fallacy discussed in the cognitive science literature (Tversky and Kahneman, 1983; Kahneman, 2011) to construct synthetic datasets on which we perform our token perturbation. This section briefly introduces the underlying logic.
Conjunction Fallacy
The most often-cited example of conjunction fallacy is called the Linda problem which is framed as follows  (Tversky and Kahneman, 1983): Linda is 31 years old, single, outspoken, and very bright. She majored in philosophy. As a student, she was deeply concerned with issues of discrimination and social justice, and also participated in antinuclear demonstrations. Which is more probable?
(a) Linda is a bank teller.
(b) Linda is a bank teller and is active in the feminist movement.
Tversky and Kahneman (1983) found that humans tend to prefer option (b). However, it is logically necessary that the probability of a conjunction of two events (e.g., Linda is a bank teller, and she is active in the feminist movement) is less than the probability of either event alone.
Syllogistic Fallacy
The syllogistic fallacy documents the logical failure that occurs when people are presented with syllogisms – two premises followed by a conclusion. Ideally, if the premises are true and the logical structure is valid, the conclusion must necessarily be true. However, when the argument’s structure is flawed, the conclusion may be invalid despite the surface-level logical form. Consider the following syllogism from Kahneman (2011): Is this logically sound?
All roses are flowers.
Some flowers fade quickly.
Therefore some roses fade quickly.
The argument is incorrect because the two premises do not imply that the set of roses and the set of flowers that fade quickly necessarily overlap.
3.2 Lost in Irrelevant Context
Logical fallacies often contain misleading contexts, exploiting common cognitive biases and shortcuts in human reasoning. These fallacies can seem convincing at first glance, being effective in swaying opinions, because they resonate with intuitive yet flawed biases. For instance, conjunction fallacies present two options: one involving a single event and the other with an additional event in conjunction. This added event is particularly designed to align with the contextual background in the problem statement, leading humans or LLMs to reaffirm their preexisting beliefs. In contrast, when the additional event in the options is changed to an irrelevant one, the model is less likely to be distracted by these extraneous and irrelevant details.
  Hypothesis 1 Genuine reasoning LLMs should withstand contextually misleading options in the problem statements.
Token perturbation: Assume problem $P$ is a conjunction fallacy problem with options $(a)$ and $(b)$. One option contains a event $x$ and the other contains $x$ and $y$ in conjunction. $y$ is relevant to the context of the problem statement that might mislead the LLM. In contrast, the perturbed problem $P^{\prime}$ replaces $y$ with a randomly generated $y^{\prime}$ irrelevant to the context.
$\boldsymbol{H_{0}}$: $\pi_{12}=\pi_{21}$.
$\boldsymbol{H_{a}}$: $\pi_{12}<\pi_{21}$. ($\pi_{12}>\pi_{21}$ is invalid.)
Here is an example of such token perturbations, represented by the right arrow mark: Kai is a community leader of Pacific Islander descent. He holds degrees in Public Administration and is passionate about preserving his cultural heritage. Which is more probable?
(a) Kai is a law enforcement worker.
(b) Kai is a law enforcement worker and participates in cultural preservation organizations $\rightarrow$ learns to play the ukulele.
To further explain $\boldsymbol{H_{0}}$ and $\boldsymbol{H_{a}}$, the null hypothesis $\boldsymbol{H_{0}}$ always assumes that the LLM is a genuine reasoner and can consistently perform reasoning regardless of the superficial token changes, leading us to expect $\pi_{12}=\pi_{21}$. Meanwhile, token bias can systematically and predictably influence the LLM’s performance. Accordingly, $\pi_{12}<\pi_{21}$ aligns with our expectation that misleading tokens degrade performance, but an observation of $\pi_{12}>\pi_{21}$ where non-misleading options result in a greater decrease in performance - would be considered invalid.
3.3 Token Bias on Widely Cited Examples in Classic Literature
[…]
  Hypothesis 2 Genuine reasoning LLM should withstand surface-level alterations to the one-shot exemplar in the problem statements.
Token perturbation: Assume one-shot in-context learning scenarios. $P$ has the original Linda problem as the one-shot exemplar. In contrast, the perturbed problem $P^{\prime}$ rephrases the exemplar to a persona called "Bob".
$\boldsymbol{H_{0}}$: $\pi_{12}=\pi_{21}$.
$\boldsymbol{H_{a}}$: $\pi_{12}>\pi_{21}$. ($\pi_{12}<\pi_{21}$ is invalid.)
3.4 Token Bias on Well-Known Entity Names
Celebrity names inherently carry a rich contextual background that LLMs learn from massive training data. We hypothesize that by replacing a celebrity name with a generic one in a conjunction fallacy problem, thereby dissociating the link to this contextual backdrop, we might see performance improvements in LLMs, and such results would underscore the potential deficiency in their genuine reasoning capabilities.
  Hypothesis 3 Genuine reasoning LLMs should withstand irrelevant alterations to name entities in problem statements
Token perturbation: Assume $P$ is a conjunction fallacy problem that involves a celebrity. In contrast, the perturbed problem $P^{\prime}$ replaces the celebrity name with a generic one.
$\boldsymbol{H_{0}}$: $\pi_{12}=\pi_{21}$.
$\boldsymbol{H_{a}}$: $\pi_{12}<\pi_{21}$. ($\pi_{12}>\pi_{21}$ is invalid.)
Token perturbation: Assume $P$ is a mathematical reasoning problem that involves a classic entity name in its story narratives. In contrast, the perturbed problem $P^{\prime}$ alters the classic name to another random one.
$\boldsymbol{H_{0}}$: $\pi_{12}=\pi_{21}$.
$\boldsymbol{H_{a}}$: $\pi_{12}>\pi_{21}$. ($\pi_{12}<\pi_{21}$ is invalid.)
[…]
(a) Her first show is a flop.
(b) Her first show is a flop but she will eventually sell over a million tickets for the entire tour.
Here is another example of the token perturbation applied to the classic "twenty-five horses" problem in mathematical reasoning, as referenced in Figure 2. Note that perturbing the numbers is optional, but the total number of animals must always be a square number: There are $\rightarrow$ thirty-six $\rightarrow$ bunnies among which you need to find out the fastest three. You can conduct a race among at most $\rightarrow$ six to find out their relative speed. At no point, you can find out the actual speed of the $\rightarrow$ bunnies in a race. Find out the minimum number of races which are required to get the top $\rightarrow$ six $\rightarrow$ bunnies.
3.5 Token Bias in Reasoning about Sets
The syllogistic fallacy involves reasoning about sets, utilizing specific quantifiers such as "all" and "some" to specify the distribution of variables. Our investigation centers on whether LLMs overfit to these tokens of quantifiers, relying heavily on their presence to generate answers that appear correct. By rephrasing these tokens with other words that convey the same meaning, we can test the robustness of LLMs’ reasoning abilities.
  Hypothesis 4 Genuine reasoning LLM should withstand irrelevant alterations to the quantifiers in problem statements.
Token perturbation: Assume $P$ is a syllogistic fallacy problem with quantifier tokens like "All" and "Some". In contrast, the perturbed problem removes "All" or rephrases "all" and "some" to different words with the same meaning.
$\boldsymbol{H_{0}}$: $\pi_{12}=\pi_{21}$.
$\boldsymbol{H_{a}}$: $\pi_{12}>\pi_{21}$. ($\pi_{12}<\pi_{21}$ is invalid.)
[…]
Continuing with the exploration of token bias in syllogistic fallacies, we propose an intriguing rephrasing of the syllogism’s narrative by incorporating the names of reputable news agencies and universities. While adding the tokens of their names does not alter the logic, it could influence how LLMs perceive and process the information. LLMs prone to token bias might erroneously increase their confidence in the trustworthiness and credibility of the stories, based purely on the association with these respected institutions.
  Hypothesis 5 Genuine reasoning LLM should withstand alterations to the narrative.
Token perturbation: Assume $P$ is the original problem. The perturbed problem $P^{\prime}$ adds or modifies specific tokens in the problem statement to reframe its narratives without changing the logic structure.
$\boldsymbol{H_{0}}$: $\pi_{12}=\pi_{21}$.
$\boldsymbol{H_{a}}$: $\pi_{12}<\pi_{21}$ or $\pi_{12}>\pi_{21}$.
[…]
3.6 Leaking Hint Tokens
Just as a proficient student doesn’t need hints to excel in a math exam, a reasoning agent should solve logical problems effectively without explicit cues. Besides, even if a student answers all problems correctly but the examlet provides all the reasoning steps, we may still question whether the student really understands the reasoning. Our experiments deliberately leak important hints that we expect a genuine reasoner to figure out itself in its intermediate reasoning steps.
  Hypothesis 6 Genuine Reasoning LLMs should not rely on hint tokens to derive correct inferences.
Token perturbation: Assume in-context learning scenarios. The perturbed problem $P^{\prime}$ explicitly adds hint tokens in its prompts, such as the name of the logical fallacy or detailed guidance on the correct reasoning, while $P$ does not.
$\boldsymbol{H_{0}}$: $\pi_{12}=\pi_{21}$.
$\boldsymbol{H_{a}}$: $\pi_{12}<\pi_{21}$. ($\pi_{12}>\pi_{21}$ is invalid.)
Here is an example of such token perturbations: Marsha Ellis, 42, is an African American transgender female. She is an ardent advocate for gender-affirming rights and environmental protection. Which is more probable?$\rightarrow$ Please be aware that this is a problem on the conjunction fallacy.
(a) Marsha is a research scientist.
(b) Marsha is a research scientist and volunteers at LGBTQ+ health centers.
We also manually craft a detailed chain-of-thought instructions Wei et al. (2022b) that teach LLMs about correct reasoning steps and potential logical pitfalls, as shown in Appendix B.
4 Experiment
Our experiments aim to rigorously test the reasoning capabilities of LLMs through the hypotheses in Section 3 on token bias. More comprehensive results are included in Appendix D. In all experiments, we run $n$ trials for each "model-prompting method" pair, depending on how many synthetic data samples are related to each hypothesis, and then perform a McNemar test. We apply the Benjamini-Hochberg Procedure and reject the null hypothesis if the p-value is less than $\alpha=0.05$.
4.1 Models
We experiment with a variety of commercial and open-sourced LLMs for a thorough study, including OpenAI gpt-3.5-turbo, gpt-4-turbo, and gpt-4o Brown et al. (2020); Achiam et al. (2023), Meta llama-3-70b-instruct, llama-3-8b-instruct, and llama-2-70b-chat Touvron et al. (2023), Anthropic claude-3-opus-20240229 and claude-3-sonnet-20240229 Anthropic (2024), and Mistral mistral-large-largest Jiang et al. (2023).
4.2 Synthetic Dataset Generation
We leverage data sources such as occupational statistics (USDL, 2024), commonsense stories (Mostafazadeh et al., 2016), CNN news stories (See et al., 2017), common disease symptom pairs kag , celebrity names Rosenberg (2021); Wikipedia contributors (2024a), objects vocabularies esl and common U.S. news media Wikipedia contributors (2024b); Pew Research Center (2011) to curate lists of entities to generate synthetic data. We outline our well-controlled data generation process and the samples used for each hypothesis in Appendix C.
(a) Experimental results for Hypothesis 3.2 ($n=400$). The perturbed problems alternate options contextually relevant to the problem statements to irrelevant ones. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs fail to reason against contextually misleading options in conjunction fallacy problems.
(b) Experimental results for Hypothesis 3.3 ($n=500$). The perturbed problems alternate the name classic "Linda" to "Bob" in in-context learning exemplars. We run one-shot with and without chain-of-thought prompts. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}$. We conclude that LLMs possess strong token bias to the name "Linda" frequently appearing in classic literature.
(c) Experimental results for Hypothesis with celebrity names 3.4 ($n=100$). The perturbed problems alternate the celebrity name to a generic one in problem statements. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs are frequently misled by celebrity names in problem statements that are irrelevant to logical essence.
(d) Experimental results for Hypothesis 3.5 ($n=200$). The perturbed problems alternate tokens "All" and "Some" to different but equivalent expressions in syllogisms. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that most LLMs rely on patterns "All…, Some…, Some…" for reasoning about syllogism.
(e) Experimental results for Hypothesis 3.5 ($n=200$). The perturbed problems add the names of trustworthy news agencies and universities to alter the narratives of syllogisms. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs might be misled by reputable names irrelevant to the logical structure.
(f) Experimental results for Hypothesis 3.6 ($n=800$). The perturbed problems leak hint tokens, either weak or strong hints in problem statements. We run zero-shot and one-shot prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs still heavily rely on hint tokens for solving logical fallacy problems well.
Figure 4: Our controlled experiments cast doubt on the genuine reasoning capabilities of LLMs. In this figure, each pair of histograms stuck together represents a comparison in the contingency table 1 for McNemar’s Tests.
4.3 Prompting Methods
We implemented commonly used prompting strategies that are sufficient for evaluating the null hypotheses within our framework. The specific prompting techniques we utilized are as follows, with their corresponding notations presented in Figure 4: Baseline: Directly answering the question without additional instructions. Zero-shot chain-of-thought (zs_cot): Includes the instruction "Let us think step by step" Wei et al. (2022b). One-shot (os): Involves a single in-context learning example Brown et al. (2020). Few-shots (fs): Utilizes three in-context examples. Similarly, we have os_cot and fs_cot. We also include weak_control_zs/os_cot and control_zs/os_cot for Hypothesis 3.6 representing prompts with additional weak or strong hints, as detailed in Appendix B.
4.4 Hypothesis Testing Results
Testing of Hypothesis 1: LLMs Would Fail at Misleading Options
We evaluate LLMs on all conjunction fallacy problems with misleading options ($n$=400). Figure 4(a) and Table LABEL:table:h1 show a significant decline in success rate when contextually misleading options in conjunction fallacy problems are replaced with random alternatives. The random ones are no longer relevant to the problem statements, so all LLMs become less likely to be swayed by background information that is not logically important. We, therefore, reject almost all null hypotheses.
Testing of Hypothesis 2: LLMs Would Fail Due to Surface Level Change in the Exemplar
We evaluate LLMs under in-context learning scenarios for solving conjunction fallacies ($n$=500). Figure 4(b) and Table LABEL:table:h2 show consistent performance drop on all LLMs when the name "Linda," frequently used in classic reasoning tasks, is substituted with "Bob" in one-shot exemplars. Such a change should not influence outcomes for genuine reasoners, as the specific name used is irrelevant to the logical process.
Testing of Hypothesis 3: LLMs Would be Misled by Celebrity Names
We evaluate LLMs on variants of conjunction fallacies that contain a celebrity name ($n$=100). We observe in Figure 4(c) and Table LABEL:table:h3 to Figure 4(a) that celebrity names appeared in problem statements frequently mislead LLMs into the celebrity’s background, which is not helpful in solving logical fallacy problems but reduces accuracy, leading us to reject all null hypotheses.
Testing of Hypothesis 4: LLMs Would Fail at Synonyms of Classic Quantifiers
We assess LLMs on syllogisms ($n$=200). Figure 4(d) and Table LABEL:table:h4 reveal that, in most instances, we should reject the null hypotheses, except for llama-3-70b-instruct. Most LLMs demonstrate insufficient robustness when patterns "All…, Some…, Some…" commonly used in classic syllogistic fallacy problems are substituted with synonyms.
Testing of Hypothesis 5: LLMs Would Be Misled by Names Linked to Reputable Entities
We evaluated the impact of names linked to reputable data sources in syllogisms ($n$=200). Figure 4(e) and Table LABEL:table:h5 demonstrate that some LLMs are indeed misled by the inclusion of these authoritative names, especially GPT-4 and LLaMA-3-70B. Generally, LLMs tend to falsely believe that these narratives are more trustworthy and, thereby, ignore the logical fallacy in them. As a result, we reject about half of the null hypotheses. Results from using the names of less credible sources are shown in Appendix LABEL:table:h5_framing for comparison.
Testing of Hypothesis 6: LLMs Still Heavily Rely on Hint Tokens
We evaluated the performance of LLMs with and without the presence of hints ($n$=200). Figure 4(f) and Table LABEL:table:h6 indicate that LLMs still heavily rely on hints to achieve ideal performance, so we reject all the null hypotheses.
5 Related Work
The Reasoning Capabilities of LLMs
There is an increasing number of works aim to improving the reasoning performance of LLMs Ouyang et al. (2022); Zhou et al. (2022); Lyu et al. (2023); Hao et al. (2023); Xu et al. (2024); Putta et al. (2024); Kumar et al. (2024); Yao et al. (2024); Cai et al. (2024), evaluating and critiquing their reasoning abilities Zhou et al. (2020); Hong et al. (2023); Huang et al. (2023); Shi et al. (2023); Sprague et al. (2024); Turpin et al. (2024); Xiao et al. (2024), trying to understand their reasoning processes Wei et al. (2022a); Merrill and Sabharwal (2023); Ma et al. (2023); Lanham et al. (2023); Merrill and Sabharwal (2024); Yang et al. (2024); Chen et al. (2024); Jin et al. (2024), as well as related surveys Qiao et al. (2022); Huang and Chang (2022); Ahn et al. (2024); Giadikiaroglou et al. (2024); Liang et al. (2024); Zheng et al. (2024) that comprehensively explore the reasoning capabilities of LLMs. Our work aligns with efforts that critique the generalization of LLMs’ reasoning capabilities, moving beyond accuracy-based benchmarks Talmor et al. (2018); Srivastava et al. (2022a, b); Fu et al. (2023); Yang et al. (2018), which primarily focus on overall question-answering accuracy, often without considering question augmentations or variations. Instead, we step one-level deeper into these problem statements, investigating potential token biases that may create the illusion of improved logical reasoning performance, while remaining vulnerable to irrelevant token perturbations. By examining these subtleties, our approach provides a more nuanced understanding of the limitations in LLM’s reasoning capabilities. Besides, we reformulate the evaluation as hypothesis testing, providing results with statistical guarantees.
Cognitive Biases and Logical Fallacies in LLMs
Recent studies Gardner et al. (2020); Hagendorff et al. (2023); Lin and Ng (2023); Talboy and Fuller (2023); Binz and Schulz (2023); Ullman (2023); Mitchell and Krakauer (2023) analyze the biases in LLMs with synthetic datasets. For example, Tamkin et al. (2023) uses an LLM to generate prompts that reveals patterns of discriminations in LLMs. Echterhoff et al. (2024) proposes a set of LLM-simulated experiments in a specific context. Although existing works Mukherjee and Chang (2024); Macmillan-Scott and Musolesi (2024); Wang et al. (2024); Suri et al. (2024); Jin et al. (2022) study more kinds of fallacy types in human psychology, they approach problems at a coarse level and only emphasize accuracy. Our study goes into a more fine-grained level with a series of hypotheses. We provide statistical guarantees and quantitative analyses of token bias that can be carefully and systematically tuned. Besides, Gou et al. (2023) presents the Rationality of Thought (RoT), decomposing responses into six predefined steps with hand-crafted prompt engineerings. Our work focuses on general prompting strategies that are sufficient to validate or reject our hypotheses.
6 Discussion
This work reconceptualizes the evaluation of the reasoning behavior of LLMs through the lens of token bias. The statistical evidence presented in our hypothesis-testing framework contributes to the larger discussion that LLMs do not always apply reasoning consistently in their decision-making processes. Instead, they primarily rely on token bias for response generation. This suggests that chain-of-thought prompting Wei et al. (2022b); Wang et al. (2022) or in-context learning Brown et al. (2020); Min et al. (2022); Lyu et al. (2022); Wang et al. (2022) may not elicit actual reasoning but instead result in semantic shortcuts for LLMs to imitate the desired behavior at superficial levels. These findings raise concerns about the extent to which LLMs truly engage in reasoning. Further investigations are needed to uncover the underlying mechanisms and limitations of LLMs’ reasoning capabilities.
7 Limitations
This hypothesis-testing framework is specifically designed for multiple-choice or yes/no questions and is not applicable to open-ended responses. It relies on LLMs with strong instruction-following capabilities to consistently produce responses that include the selected options, though we find that LLMs can generally follow these instructions in most cases. In addition, smaller LLMs, such as llama-3-8b-instruct, with lower instruction following capabilities may contain more confounders besides the token bias, which could weaken our hypothesis testing results. As a result, we mainly focus on state-of-the-art LLMs. Moreover, the finding of token biases requires manual efforts. We also acknowledge that there are likely other hypotheses and assumptions that a genuine reasoner should satisfy. Our current study focuses on the conjunction fallacy, syllogistic fallacy, the "twenty-five horses" problem in graph theory, and their variants to demonstrate our framework. These are quintessential examples and the framework could include a broader range of hypotheses, fallacy types, data modalities, and reasoning tasks.
8 Acknowledgement
This work was supported by NSF grant CCF-2112665 (TILOS), which provides funding for B.J. and C.J. It was also funded in part by the Laboratory Directed Research and Development (LDRD) program at Argonne National Laboratory, with support from the Office of Science, U.S. Department of Energy, under Contract No. DEAC02-06CH11357. Y. X. and W. J. Su acknowledge support from the NSF HDR TRIPODS award (CCF-1934876). The authors thank John K. Hutchison for his valuable suggestions regarding the motivation of this study.
Appendix A The Original Linda Problem in Tversky and Kahneman (1983)
The original Linda problem is framed as follows  (Tversky and Kahneman, 1983):
Linda is 31 years old, single, outspoken, and very bright. She majored in philosophy. As a student, she was deeply concerned with issues of discrimination and social justice, and also participated in antinuclear demonstrations. Which is more probable?
  1. 1.
    Linda is a bank teller.
  2. 2.
    Linda is a bank teller and is active in the feminist movement.
Here is an example of GPT-4o explaining the Linda Problem: https://chatgpt.com/share/eff10b9d-d219-4806-9cb9-d2d9104c0e83.
Our “Bob” version of this problem is as follows:
Bob is 29 years old, deeply passionate about environmental conservation, and volunteers his weekends at local park clean-ups. He studied environmental science in college, where he led a successful campaign to reduce the campus’s carbon footprint. Bob is also an avid cyclist and promotes sustainable living practices whenever possible. Based on this information, which is more possible?
  1. 1.
    Bob works for a renewable energy company and is an active member of a local environmental advocacy group.
  2. 2.
    Bob works for a renewable energy company.
Appendix B Prompts in Hypothesis 3.6
This section includes the detailed prompts we use to evaluate the influences from weak and strong hints. These prompts are added to either the zero-shot chain-of-thought or the one-shot chain-of-thought prompts.
B.1 Weak Hint
For Problems on Conjunction Fallacies
Your task is to answer the following question by explicitly selecting either option (a), (b), etc. Please be aware that this is a Linda Problem designed to explore the concept of the conjunction fallacy. Here is the question and let’s think step by step.
For Problems on Syllogistic Fallacies
Your task is to answer the following question by explicitly saying ’Yes’ or ’No’. Please be aware that this is a Linda Problem designed to explore the concept of the syllogistic fallacy.
B.2 Strong Hint
For Problems on Conjunction Fallacies
Your task is to answer the following question by explicitly selecting either option (a), (b), etc. Please aware that this is a Linda Problem designed to explore the concept of the conjunction fallacy. The conjunction fallacy occurs when individuals incorrectly judge the conjunction of two events as more probable than one of the events alone. For instance, many might believe that Linda, who is described as a bright, single woman deeply concerned with discrimination and social justice, is more likely to be both a bank teller and active in the feminist movement than just a bank teller. This judgment violates the basic probability rule: the probability of a conjunction, P(A and B), is always less than or equal to the probabilities of its constituents, P(A) or P(B). This error often stems from the representativeness heuristic, where people estimate the likelihood of an event by how closely it matches their mental prototype. To correctly solve problems like this, you must adopt probabilistic thinking: abstract the problem from its narrative context and focus solely on the probabilistic models. Ignore all extraneous background information and consistently choose the option involving a single event as it statistically holds a higher likelihood than the conjunction of multiple events. Here is the question and let’s think step by step.
For Problems on Syllogistic Fallacies
Your task is to answer the following question by explicitly saying ’Yes’ or ’No’. Please aware that this is a Syllogistic Fallacy Problem. This type of reasoning is known as a syllogism. Pay close attention to quantifiers such as ’All’, ’Some’, ’No’, or similar terms. These terms help define the distribution of properties or elements within the given groups or categories in the premises. Next, assess whether the attribute ascribed in the conclusion necessarily follows from the attributes described in the premises. Consider if the subset described in the second premise encompasses or overlaps with the elements in the first premise that are carried into the conclusion. A common pitfall in syllogistic reasoning is the erroneous assumption that a characteristic of a subset of a group (from the premises) applies to another subset of the same or different group (in the conclusion), without explicit justification. Ignore the background information about the objects and focus on the logical structure of the argument. Here is an example.
Appendix C Synthetic Data Generation
In this section, we outline the controlled synthetic data generation process. For each variant, we generate 100 synthetic data samples.
C.1 Conjunction Fallacy
We create several variants of the conjunction fallacy problem discussed in the original work by Tversky and Kahneman (1983):
Variant 1
The original Linda Problem. We maintain the narrative structure of the original Linda Problem described in Appendix A. We ask GPT-4 to randomly pick reasonable personal details such as name, race, gender identity, age, and major, forming a short biography. GPT-4 then crafts two options $(a)$ and $(b)$ for each problem, both of which contain the same randomly selected occupation from USDL (2024) like "Linda is a bank teller". The longer option also contains a hobby that must be relevant to the bio like "active in the feminist movement".
The prompt used to generate the bio is as follows, where {random_gender}, {random_race},{random_age} are sampled from a pre-defined random function:
Your task is to write a short bio for a random person within 100 words. You shall pick a random name, use gender {random_gender}, race {random_race}, and an age {random_age}. The bio should describe the college majors, some personal characters, and interests. Keep the bio short. For example, ’Linda is 31 years old, single, outspoken, and very bright. She majored in philosophy. As a student, she was deeply concerned with issues of discrimination and social justice, and also participated in anti-nuclear demonstrations. Write another example here:
We then follow-up the conversation with the following prompt:
Your next step is to find a hobby or activity that the person mentioned before will be interested in based on your experience. The hobby or activity must be relevant to the bio descriptions. In the example above, we can say that ’Linda is active in the feminist movement.’ because her bio says she was concerned with discrimination and social justice. Please keep your answer in one sentence and begin with that person’s name, but refrain from using any words used in the bio.
To create token bias, we generate a random hobby using the following:
Your task is to find a random hobby or activity, and keep your answer short in one sentence. For example, you can say ’cook Asian foods.’
Variant 2
In the original paper, the following variant of the conjunction fallacy problem is also presented:
John P. is a meek man, 42 years old, married with two children. His neighbors describe him as mild-mannered but somewhat secretive. He owns an import-export company based in New York City, and he travels frequently to Europe and the Far East. Mr. P. was convicted once for smuggling precious stones and metals (including uranium) and received a suspended sentence of 6 months in jail and a large fine. Mr. P. is currently under police investigation. Which one is more likely?
  1. 1.
    Mr. P. killed one of his employees.
  2. 2.
    Mr. P. killed one of his employees to prevent him from talking to the police.
The conjunction of two events in the second option is connected by the word ‘to.’ To create this dataset, we sample a random story from the collection of commonsense stories (Mostafazadeh et al., 2016) and CNN news stories (See et al., 2017). We use all the sentences in the story as the context of the conjunction fallacy problem except for the last one. We use the last sentence as the first option in the problem. As for the second option, we append the last sentence, add the connecting word ‘to,’ and then we prompt GPT-4 to complete the second option. The prompt used here is similar to that discussed in section 2.
To perform token perturbation, we further prompt GPT-4 with the following:
Your next task is to complete the last sentence of the same problem but make sure your completion after ’to’ is now irrelevant to the content intentionally:
Variant 3
This is the same as the last variant, except that we use ‘because’ as the connecting word.
Variant 4
This is the same as the last variant, except that we use ‘so that’ as the connecting word.
Variant 5
In the original paper, the following variant of the conjunction fallacy is discussed:
A 55-year-old woman had pulmonary embolism documented angiographically 10 days after a cholecystectomy. Which is more likely?
  1. 1.
    dyspnea and hemiparesis
  2. 2.
    hemiparesis
Inspired by this example, we randomly sample a disease and its corresponding symptoms from kag and apply the following prompt to generate a conjunction fallacy problem:
Your task is to create another conjunction fallacy quiz following the format in the example below. Do not mention the name ’conjunction fallacy.’ You should pick a random name for the patient, use gender {random_gender} race {random_race}, an age {random_age} and the disease {random_disease} in your new problem statement. The question should be ’Which one is more likely?’ followed by two options (a) and (b), one of which should be a subset of the other. You can randomly switch the order of which option is (a) and which is (b). You should use the symptoms {random_symptom_one} in both options and add {random_symptom_two} to the longer option only. Do not make any changes to the given disease or the symptoms. Here is the new problem:
We then prompt GPT-4 for an irrelevant symptom:
Your task is to create another conjunction fallacy quiz following the format in the example below. Do not mention the name ’conjunction fallacy.’ You should pick a random name for the patient, use gender {random_gender} race {random_race}, an age {random_age} and the disease {random_disease} in your new problem statement. The question should be ’Which one is more likely?’ followed by two options (a) and (b), one of which should be a subset of the other. You can randomly switch the order of which option is (a) and which is (b). You should use the symptoms {random_symptom_one} in both options. You should add another random symptoms to the longer option only, which must be completely irrelevant to the disease {random_disease} intentionally. Do not make any changes to the given disease or the symptoms. Here is the new problem:
Variant 6
In the original paper, the following variant of the conjunction fallacy is discussed:
Suppose Bjorn Borg reaches the Wimbledon finals in 1981. Which is more likely?
  1. 1.
    Borg will lose the first set
  2. 2.
    Borg will lose the first set but win the match
Inspired by this example, we randomly sample celebrity names from the Times Person of the Year Rosenberg (2021) and Forbes Celebrity 100 Wikipedia contributors (2024a) and apply the following few-shot prompt to generate a conjunction fallacy problem.
Create one example that look like this:
Suppose [celebrity is going to do something]. Which is more likely:
(a) [Something unlikely for this person]
(b) [Something unlikely for this person] but [something extremely likely for this person]
Here are some examples:
Suppose Taylor Swift is going to have another tour in 2027. Which is more likely:
(a) Her first show is a flop.
(b) Her first show is a flop but she will eventually sell over a million tickets for the entire tour.
Suppose Barack Obama is running for president in 2024. Which is more likely:
(a) Barack Obama will win the national popular vote
(b) Barack Obama will win the national popular vote but lose the Electoral College vote
Suppose Bjorn Borg reaches the Wimbledon finals. Which outcome is more likely?
(a) Borg will lose the first set
(b) Borg will lose the first set but win the match
Complete the following. Do not output anything else.
Suppose {random_celebrity}
For Hypothesis 3.2, we include Variant 2,3,4 and 5, resulting in $n=400$ samples. For Hypothesis 3.3, we include Variant 2,3,4,5 and 6, resulting in $n=500$ samples. For Hypothesis 3.4, we include Variant 5, resulting in $n=100$ samples.
C.2 Syllogistic Fallacy
For Hypothesis 3.5, we randomly sample an entity {random_object} from a curated list of objects from esl and use the following few-shot prompt to generate $n=200$ problems:
Fill in the blanks in the following template. Do not output anything else.
All [objects] are [category].
Some [category]s [characteristic traits of this category].
Therefore some [same objects as before] [characteristic traits this category].
Make sure that the characteristic traits of this category only fit for a subset of this category but not for all.
For example:
All carrots are vegetables.
Some vegetables are rich in fiber.
Therefore, some carrots are rich in fiber.
All roses are flowers.
Some flowers fade quickly.
Therefore some roses fade quickly.
All actors are performers.
Some performers are skilled in improvisation.
Therefore some actors are skilled in improvisation.
All {random_object} are
The common U.S. news media sources we used to perturb these problems in Hypothesis 3.5 are taken from Wikipedia contributors (2024b); Pew Research Center (2011). This also results in $n=200$ samples.
Appendix D Additional Experiment Results
D.1 Hypothesis 1
The full experimental results for Hypothesis 3.2 are shown in Figure 5, 6 and Table LABEL:table:h1.
Table 2: Full Experimental results for Hypothesis 3.2. Note that in our experiments, $n^{*}=n_{21}+n_{12}$ is not equal to the number of data samples $n$. Here, $n_{12}$ denotes the instances where the LLM correctly answers the original problem but fails on the perturbed version, and $n_{21}$ denotes the opposite scenario. Thus, a large value of $n^{*}$ happens only when the LLM makes many mistakes. Specifically, GPT-4o in this table shows $n^{*}=1$ with few-shots learning. We find that GPT-4o is excellent in answering these problems with few-show exemplars, achieving near-perfect accuracy of almost $100\%$, so their $n_{12}$ and $n_{21}$ values are pretty small. This high accuracy, however, only lead us to fail to reject this particular instantiation of the null hypothesis, but not in other situations. While we could increase the sample size from the current $n$ to potentialy observe more errors, thus a higher $n^{*}$ in scenarios involving state-of-the-art LMs with few-shot learning, our rejection of the null hypothesis under other scenarios when tested against the GPT-4o is sufficient to argue that LLMs are not yet genuine reasoners.
model prompting method $n_{12}$ $n_{21}$ $n^{*}$ z-stat p-value reject
gpt-3.5-turbo baseline 4 160 164 12.181553 0.000000 True
gpt-3.5-turbo zs-cot 19 218 237 12.926439 0.000000 True
gpt-3.5-turbo os 3 115 118 10.310436 0.000000 True
gpt-3.5-turbo os-cot 17 147 164 10.151295 0.000000 True
gpt-3.5-turbo fs 12 132 144 10.000000 0.000000 True
gpt-3.5-turbo fs-cot 6 180 186 12.758299 0.000000 True
gpt-4-turbo baseline 7 357 364 18.344985 0.000000 True
gpt-4-turbo zs-cot 6 331 337 17.703878 0.000000 True
gpt-4-turbo os 0 73 73 8.544004 0.000000 True
gpt-4-turbo os-cot 1 101 102 9.901475 0.000000 True
gpt-4-turbo fs 1 38 39 5.924742 0.000000 True
gpt-4-turbo fs-cot 0 50 50 7.071068 0.000000 True
gpt-4o baseline 5 360 365 18.581549 0.000000 True
gpt-4o zs-cot 3 281 284 16.496265 0.000000 True
gpt-4o os 0 33 33 5.744563 0.000000 True
gpt-4o os-cot 8 101 109 8.907784 0.000000 True
gpt-4o fs 0 1 1 1.000000 0.158655 False
gpt-4o fs-cot 3 72 75 7.967434 0.000000 True
llama-2-70b-chat baseline 3 215 218 14.358452 0.000000 True
llama-2-70b-chat zs-cot 30 199 229 11.167834 0.000000 True
llama-2-70b-chat os 56 101 157 3.591391 0.000181 True
llama-2-70b-chat os-cot 38 170 208 9.152553 0.000000 True
llama-2-70b-chat fs 90 115 205 1.746076 0.042775 True
llama-2-70b-chat fs-cot 43 150 193 7.702029 0.000000 True
meta-llama-3-70b-instruct baseline 10 321 331 17.094106 0.000000 True
meta-llama-3-70b-instruct zs-cot 12 296 308 16.182402 0.000000 True
meta-llama-3-70b-instruct os 22 32 54 1.360828 0.090122 False
meta-llama-3-70b-instruct os-cot 22 58 80 4.024922 0.000032 True
meta-llama-3-70b-instruct fs 18 26 44 1.206045 0.116049 False
meta-llama-3-70b-instruct fs-cot 19 44 63 3.149704 0.000883 True
meta-llama-3-8b-instruct baseline 8 272 280 15.777018 0.000000 True
meta-llama-3-8b-instruct zs-cot 5 263 268 15.759858 0.000000 True
meta-llama-3-8b-instruct os 12 102 114 8.429272 0.000000 True
meta-llama-3-8b-instruct os-cot 19 154 173 10.263860 0.000000 True
meta-llama-3-8b-instruct fs 26 77 103 5.025179 0.000000 True
meta-llama-3-8b-instruct fs-cot 20 112 132 8.007572 0.000000 True
claude-3-opus-20240229 baseline 17 241 258 13.945631 0.000000 True
claude-3-opus-20240229 zs-cot 10 229 239 14.165932 0.000000 True
claude-3-opus-20240229 os 15 136 151 9.846840 0.000000 True
claude-3-opus-20240229 os-cot 13 157 170 11.044296 0.000000 True
claude-3-opus-20240229 fs 9 140 149 10.731938 0.000000 True
claude-3-opus-20240229 fs-cot 14 129 143 9.616783 0.000000 True
claude-3-sonnet-20240229 baseline 8 364 372 18.457740 0.000000 True
claude-3-sonnet-20240229 zs-cot 8 313 321 17.023440 0.000000 True
claude-3-sonnet-20240229 os 0 258 258 16.062378 0.000000 True
claude-3-sonnet-20240229 os-cot 9 165 174 11.826329 0.000000 True
claude-3-sonnet-20240229 fs 3 175 178 12.891945 0.000000 True
claude-3-sonnet-20240229 fs-cot 24 143 167 9.208496 0.000000 True
mistral-large-latest baseline 8 388 396 19.095718 0.000000 True
mistral-large-latest zs-cot 5 384 389 19.216063 0.000000 True
mistral-large-latest os 3 91 94 9.076507 0.000000 True
mistral-large-latest os-cot 4 79 83 8.232319 0.000000 True
mistral-large-latest fs 5 74 79 7.763107 0.000000 True
mistral-large-latest fs-cot 5 79 84 8.074062 0.000000 True
Figure 5: Full experimental results for Hypothesis 3.2 ($n=400$). The perturbed problems alternate options contextually relevant to the problem statements to irrelevant ones. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs fail to reason against contextually misleading options in conjunction fallacy problems.
Figure 6: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.2.
D.2 Hypothesis 2
The full experimental results for Hypothesis 3.3 are shown in Figure 7, 8 and Table LABEL:table:h2.
Table 3: Full experimental results for Hypothesis 3.3
model prompting method $n_{12}$ $n_{21}$ $n^{*}$ z-stat p-value reject
gpt-3.5-turbo os 164 18 182 -10.822240 0.000000 True
gpt-3.5-turbo os-cot 160 54 214 -7.246011 0.000000 True
gpt-4-turbo os 110 13 123 -8.746195 0.000000 True
gpt-4-turbo os-cot 109 12 121 -8.818182 0.000000 True
gpt-4o os 14 18 32 0.707107 0.760250 False
gpt-4o os-cot 64 22 86 -4.528976 0.000003 True
llama-2-70b-chat os 62 48 110 -1.334848 0.096314 False
llama-2-70b-chat os-cot 111 43 154 -5.479596 0.000000 True
meta-llama-3-70b-instruct os 253 18 271 -14.275233 0.000000 True
meta-llama-3-70b-instruct os-cot 243 15 258 -14.194660 0.000000 True
meta-llama-3-8b-instruct os 241 33 274 -12.565740 0.000000 True
meta-llama-3-8b-instruct os-cot 162 54 216 -7.348469 0.000000 True
claude-3-opus-20240229 os 157 73 230 -5.538796 0.000000 True
claude-3-opus-20240229 os-cot 151 72 223 -5.290231 0.000000 True
claude-3-sonnet-20240229 os 162 9 171 -11.700202 0.000000 True
claude-3-sonnet-20240229 os-cot 166 36 202 -9.146768 0.000000 True
mistral-large-latest os 286 18 304 -15.370854 0.000000 True
mistral-large-latest os-cot 275 34 309 -13.710011 0.000000 True
Figure 7: Full experimental results for Hypothesis 3.3 ($n=500$). The perturbed problems alternate the name classic "Linda" to "Bob" in in-context learning exemplars. We run one-shot with and without chain-of-thought prompts. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}$. We conclude that LLMs possess a strong token bias toward the name "Linda", which frequently appears in classic literature.
Figure 8: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.3.
D.3 Hypothesis 3
The full experimental results for Hypothesis 3.4 are shown in Figure 9, 10 and Table LABEL:table:h3.
Table 4: Full experimental results for Hypothesis 3.4
model prompting method $n_{12}$ $n_{21}$ $n^{*}$ z-stat p-value reject
gpt-3.5-turbo baseline 9 19 28 1.889822 0.072141 False
gpt-3.5-turbo zs-cot 14 23 37 1.479591 0.133569 False
gpt-3.5-turbo os 7 11 18 0.942809 0.341537 False
gpt-3.5-turbo os-cot 12 13 25 0.200000 0.516363 False
gpt-3.5-turbo fs 3 17 20 3.130495 0.005798 True
gpt-3.5-turbo fs-cot 12 20 32 1.414214 0.137003 False
gpt-4-turbo baseline 5 21 26 3.137858 0.004595 True
gpt-4-turbo zs-cot 1 19 20 4.024922 0.000270 True
gpt-4-turbo os 0 6 6 2.449490 0.046875 True
gpt-4-turbo os-cot 1 5 6 1.632993 0.178977 False
gpt-4-turbo fs 2 8 10 1.897367 0.113582 False
gpt-4-turbo fs-cot 3 7 10 1.264911 0.250845 False
gpt-4o baseline 3 19 22 3.411211 0.003300 True
gpt-4o zs-cot 1 7 8 2.121320 0.075938 False
gpt-4o os 0 1 1 1.000000 0.562500 False
gpt-4o os-cot 2 2 4 0.000000 0.727941 False
gpt-4o fs 0 0 0 0.000000 1.000000 False
gpt-4o fs-cot 2 4 6 0.816497 0.441964 False
llama-2-70b-chat baseline 7 6 13 -0.277350 0.736760 False
llama-2-70b-chat zs-cot 10 14 24 0.816497 0.361423 False
llama-2-70b-chat os 3 17 20 3.130495 0.005798 True
llama-2-70b-chat os-cot 9 20 29 2.042649 0.055468 False
llama-2-70b-chat fs 8 16 24 1.632993 0.136431 False
llama-2-70b-chat fs-cot 12 11 23 -0.208514 0.714075 False
meta-llama-3-70b-instruct baseline 4 20 24 3.265986 0.004595 True
meta-llama-3-70b-instruct zs-cot 3 27 30 4.381780 0.000106 True
meta-llama-3-70b-instruct os 3 6 9 1.000000 0.351562 False
meta-llama-3-70b-instruct os-cot 5 4 9 -0.333333 0.760171 False
meta-llama-3-70b-instruct fs 6 6 12 0.000000 0.675323 False
meta-llama-3-70b-instruct fs-cot 8 9 17 0.242536 0.562500 False
meta-llama-3-8b-instruct baseline 4 18 22 2.984810 0.009021 True
meta-llama-3-8b-instruct zs-cot 9 25 34 2.743977 0.011706 True
meta-llama-3-8b-instruct os 1 2 3 0.577350 0.562500 False
meta-llama-3-8b-instruct os-cot 9 19 28 1.889822 0.072141 False
meta-llama-3-8b-instruct fs 4 6 10 0.632456 0.473383 False
meta-llama-3-8b-instruct fs-cot 13 13 26 0.000000 0.562500 False
claude-3-opus-20240229 baseline 6 24 30 3.286335 0.003426 True
claude-3-opus-20240229 zs-cot 7 28 35 3.549648 0.001736 True
claude-3-opus-20240229 os 1 7 8 2.121320 0.075938 False
claude-3-opus-20240229 os-cot 4 7 11 0.904534 0.361423 False
claude-3-opus-20240229 fs 5 12 17 1.697749 0.133569 False
claude-3-opus-20240229 fs-cot 7 17 24 2.041241 0.075030 False
claude-3-sonnet-20240229 baseline 4 33 37 4.767571 0.000050 True
claude-3-sonnet-20240229 zs-cot 11 25 36 2.333333 0.033127 True
claude-3-sonnet-20240229 os 1 9 10 2.529822 0.034122 True
claude-3-sonnet-20240229 os-cot 5 10 15 1.290994 0.226318 False
claude-3-sonnet-20240229 fs 6 17 23 2.293659 0.049296 True
claude-3-sonnet-20240229 fs-cot 7 13 20 1.341641 0.203021 False
mistral-large-latest baseline 3 22 25 3.800000 0.000781 True
mistral-large-latest zs-cot 2 25 27 4.426352 0.000106 True
mistral-large-latest os 0 3 3 1.732051 0.198529 False
mistral-large-latest os-cot 0 4 4 2.000000 0.125000 False
mistral-large-latest fs 2 12 14 2.672612 0.023291 True
mistral-large-latest fs-cot 5 11 16 1.500000 0.177283 False
Figure 9: Full experimental results for Hypothesis 3.4 ($n=100$). The perturbed problems alternate the celebrity name to a generic one in problem statements. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs are frequently misled by irrelevant celebrity names in problem statements that are irrelevant to logical essence.
Figure 10: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.4.
D.4 Hypothesis 4
The full experimental results for Hypothesis 3.5 are shown in Figure 11, 12 and Table LABEL:table:h4.
Table 5: Full experimental results for Hypothesis 3.5
model prompting method $n_{12}$ $n_{21}$ $n^{*}$ z-stat p-value reject
gpt-3.5-turbo baseline 21 1 22 -4.264014 0.000023 True
gpt-3.5-turbo zs-cot 20 3 23 -3.544745 0.000732 True
gpt-3.5-turbo os 3 2 5 -0.447214 0.771429 False
gpt-3.5-turbo os-cot 14 13 27 -0.192450 0.714985 False
gpt-3.5-turbo fs 1 0 1 -1.000000 0.771429 False
gpt-3.5-turbo fs-cot 7 9 16 0.500000 1.000000 False
gpt-4-turbo baseline 33 0 33 -5.744563 0.000000 True
gpt-4-turbo zs-cot 60 3 63 -7.181325 0.000000 True
gpt-4-turbo os 28 0 28 -5.291503 0.000001 True
gpt-4-turbo os-cot 27 2 29 -4.642383 0.000008 True
gpt-4-turbo fs 14 5 19 -2.064742 0.068654 False
gpt-4-turbo fs-cot 14 3 17 -2.667892 0.014939 True
gpt-4o baseline 24 2 26 -4.314555 0.000031 True
gpt-4o zs-cot 32 4 36 -4.666667 0.000008 True
gpt-4o os 10 12 22 0.426401 1.000000 False
gpt-4o os-cot 4 2 6 -0.816497 0.618750 False
gpt-4o fs 9 0 9 -3.000000 0.005022 True
gpt-4o fs-cot 12 4 16 -2.000000 0.079767 False
llama-2-70b-chat baseline 0 0 0 0.000000 1.000000 False
llama-2-70b-chat zs-cot 0 0 0 0.000000 1.000000 False
llama-2-70b-chat os 14 1 15 -3.356586 0.001388 True
llama-2-70b-chat os-cot 17 4 21 -2.836833 0.008833 True
llama-2-70b-chat fs 28 5 33 -4.003786 0.000099 True
llama-2-70b-chat fs-cot 0 0 0 0.000000 1.000000 False
meta-llama-3-70b-instruct baseline 0 3 3 1.732051 1.000000 False
meta-llama-3-70b-instruct zs-cot 7 4 11 -0.904534 0.510978 False
meta-llama-3-70b-instruct os 5 15 20 2.236068 1.000000 False
meta-llama-3-70b-instruct os-cot 14 17 31 0.538816 1.000000 False
meta-llama-3-70b-instruct fs 5 13 18 1.885618 1.000000 False
meta-llama-3-70b-instruct fs-cot 13 7 20 -1.341641 0.263176 False
meta-llama-3-8b-instruct baseline 0 2 2 1.414214 1.000000 False
meta-llama-3-8b-instruct zs-cot 2 1 3 -0.577350 0.771429 False
meta-llama-3-8b-instruct os 0 5 5 2.236068 1.000000 False
meta-llama-3-8b-instruct os-cot 10 13 23 0.625543 1.000000 False
meta-llama-3-8b-instruct fs 0 0 0 0.000000 1.000000 False
meta-llama-3-8b-instruct fs-cot 0 0 0 0.000000 1.000000 False
claude-3-opus-20240229 baseline 37 6 43 -4.727456 0.000006 True
claude-3-opus-20240229 zs-cot 43 7 50 -5.091169 0.000001 True
claude-3-opus-20240229 os 28 2 30 -4.746929 0.000006 True
claude-3-opus-20240229 os-cot 28 1 29 -5.013774 0.000002 True
claude-3-opus-20240229 fs 15 0 15 -3.872983 0.000099 True
claude-3-opus-20240229 fs-cot 9 0 9 -3.000000 0.005022 True
claude-3-sonnet-20240229 baseline 0 10 10 3.162278 1.000000 False
claude-3-sonnet-20240229 zs-cot 1 20 21 4.146140 1.000000 False
claude-3-sonnet-20240229 os 0 2 2 1.414214 1.000000 False
claude-3-sonnet-20240229 os-cot 1 2 3 0.577350 1.000000 False
claude-3-sonnet-20240229 fs 6 4 10 -0.632456 0.656628 False
claude-3-sonnet-20240229 fs-cot 1 1 2 0.000000 1.000000 False
mistral-large-latest baseline 48 0 48 -6.928203 0.000000 True
mistral-large-latest zs-cot 47 0 47 -6.855655 0.000000 True
mistral-large-latest os 16 0 16 -4.000000 0.000055 True
mistral-large-latest os-cot 26 0 26 -5.099020 0.000001 True
mistral-large-latest fs 5 2 7 -1.133893 0.436942 False
mistral-large-latest fs-cot 5 0 5 -2.236068 0.068654 False
Figure 11: Full experimental results for Hypothesis 3.5 ($n=200$). The perturbed problems alternate tokens "All" and "Some" to different but equivalent expressions in syllogisms. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that most LLMs rely on patterns "All…, Some…, Some…" for reasoning about syllogism.
Figure 12: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.5.
D.5 Hypothesis 5
The full experimental results for Hypothesis 3.5 (sets-original vs. sets-framing) are shown in Figure 13, 14 and Table LABEL:table:h5.
Table 6: Full experimental results for Hypothesis 3.5(sets-original-random vs sets-original-framing-gold)
model prompting method $n_{12}$ $n_{21}$ $n^{*}$ z-stat p-value reject
gpt-3.5-turbo baseline 1 33 34 5.487955 0.000000 True
gpt-3.5-turbo zs-cot 2 49 51 6.581316 0.000000 True
gpt-3.5-turbo os 2 18 20 3.577709 0.000836 True
gpt-3.5-turbo os-cot 13 43 56 4.008919 0.000143 True
gpt-3.5-turbo fs 0 15 15 3.872983 0.000143 True
gpt-3.5-turbo fs-cot 9 19 28 1.889822 0.083532 False
gpt-4-turbo baseline 47 2 49 -6.428571 0.000000 True
gpt-4-turbo zs-cot 23 5 28 -3.401680 0.001292 True
gpt-4-turbo os 43 4 47 -5.688735 0.000000 True
gpt-4-turbo os-cot 37 7 44 -4.522670 0.000016 True
gpt-4-turbo fs 50 4 54 -6.259807 0.000000 True
gpt-4-turbo fs-cot 46 3 49 -6.142857 0.000000 True
gpt-4o baseline 63 2 65 -7.566119 0.000000 True
gpt-4o zs-cot 57 4 61 -6.785955 0.000000 True
gpt-4o os 50 4 54 -6.259807 0.000000 True
gpt-4o os-cot 53 1 54 -7.076304 0.000000 True
gpt-4o fs 46 1 47 -6.563925 0.000000 True
gpt-4o fs-cot 45 5 50 -5.656854 0.000000 True
llama-2-70b-chat baseline 0 0 0 0.000000 1.000000 False
llama-2-70b-chat zs-cot 0 0 0 0.000000 1.000000 False
llama-2-70b-chat os 8 8 16 0.000000 1.000000 False
llama-2-70b-chat os-cot 5 0 5 -2.236068 0.086538 False
llama-2-70b-chat fs 15 21 36 1.000000 0.372495 False
llama-2-70b-chat fs-cot 0 0 0 0.000000 1.000000 False
meta-llama-3-70b-instruct baseline 3 4 7 0.377964 1.000000 False
meta-llama-3-70b-instruct zs-cot 4 10 14 1.603567 0.225501 False
meta-llama-3-70b-instruct os 24 7 31 -3.053290 0.004074 True
meta-llama-3-70b-instruct os-cot 31 12 43 -2.897473 0.006553 True
meta-llama-3-70b-instruct fs 20 6 26 -2.745626 0.010192 True
meta-llama-3-70b-instruct fs-cot 15 4 19 -2.523573 0.028816 True
meta-llama-3-8b-instruct baseline 2 46 48 6.350853 0.000000 True
meta-llama-3-8b-instruct zs-cot 0 51 51 7.141428 0.000000 True
meta-llama-3-8b-instruct os 3 37 40 5.375872 0.000000 True
meta-llama-3-8b-instruct os-cot 3 51 54 6.531973 0.000000 True
meta-llama-3-8b-instruct fs 0 53 53 7.280110 0.000000 True
meta-llama-3-8b-instruct fs-cot 0 10 10 3.162278 0.003637 True
claude-3-opus-20240229 baseline 19 28 47 1.312785 0.232268 False
claude-3-opus-20240229 zs-cot 21 28 49 1.000000 0.372495 False
claude-3-opus-20240229 os 6 29 35 3.887710 0.000228 True
claude-3-opus-20240229 os-cot 5 23 28 3.401680 0.001292 True
claude-3-opus-20240229 fs 1 6 7 1.889822 0.168750 False
claude-3-opus-20240229 fs-cot 1 3 4 1.000000 0.703125 False
claude-3-sonnet-20240229 baseline 0 49 49 7.000000 0.000000 True
claude-3-sonnet-20240229 zs-cot 0 54 54 7.348469 0.000000 True
claude-3-sonnet-20240229 os 1 28 29 5.013774 0.000001 True
claude-3-sonnet-20240229 os-cot 2 13 15 2.840188 0.011730 True
claude-3-sonnet-20240229 fs 2 18 20 3.577709 0.000836 True
claude-3-sonnet-20240229 fs-cot 1 10 11 2.713602 0.018080 True
mistral-large-latest baseline 9 3 12 -1.732051 0.187709 False
mistral-large-latest zs-cot 9 3 12 -1.732051 0.187709 False
mistral-large-latest os 2 10 12 2.309401 0.056298 False
mistral-large-latest os-cot 2 13 15 2.840188 0.011730 True
mistral-large-latest fs 3 4 7 0.377964 1.000000 False
mistral-large-latest fs-cot 1 3 4 1.000000 0.703125 False
Figure 13: Full experimental results for Hypothesis 3.5, sets-original vs. sets-framing ($n=200$). The perturbed problems add the names of trustworthy news agencies and universities to alter the narratives of syllogisms. Both the original and perturbed problems have classic patterns "All…, Some…, Some…" already removed to ensure a single confounder for token bias analysis. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs might be misled by reputable names irrelevant to the logical structure.
Figure 14: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.5.
The full experimental results for Hypothesis 3.5 (framing gold vs. random) are shown in Figure 15, 16 and Table LABEL:table:h5_framing.
Table 7: Full experimental results for Hypothesis 3.5 (sets-framing gold vs. random)
model prompting method $n_{12}$ $n_{21}$ $n^{*}$ z-stat p-value reject
gpt-3.5-turbo baseline 11 29 40 2.846050 0.011382 True
gpt-3.5-turbo zs-cot 14 25 39 1.761410 0.136166 False
gpt-3.5-turbo os 5 8 13 0.832050 0.713113 False
gpt-3.5-turbo os-cot 23 26 49 0.428571 0.767760 False
gpt-3.5-turbo fs 12 13 25 0.200000 0.927346 False
gpt-3.5-turbo fs-cot 16 12 28 -0.755929 0.596799 False
gpt-4-turbo baseline 13 13 26 0.000000 1.000000 False
gpt-4-turbo zs-cot 11 8 19 -0.688247 0.760233 False
gpt-4-turbo os 13 23 36 1.666667 0.161292 False
gpt-4-turbo os-cot 15 28 43 1.982481 0.092843 False
gpt-4-turbo fs 10 30 40 3.162278 0.004696 True
gpt-4-turbo fs-cot 10 31 41 3.279649 0.003609 True
gpt-4o baseline 4 13 17 2.182821 0.092843 False
gpt-4o zs-cot 7 13 20 1.341641 0.384095 False
gpt-4o os 6 38 44 4.824182 0.000015 True
gpt-4o os-cot 9 36 45 4.024922 0.000337 True
gpt-4o fs 2 33 35 5.239956 0.000002 True
gpt-4o fs-cot 14 26 40 1.897367 0.104003 False
llama-2-70b-chat baseline 0 19 19 4.358899 0.000029 True
llama-2-70b-chat zs-cot 0 10 10 3.162278 0.005551 True
llama-2-70b-chat os 1 74 75 8.429314 0.000000 True
llama-2-70b-chat os-cot 0 2 2 1.414214 0.637718 False
llama-2-70b-chat fs 1 44 45 6.410062 0.000000 True
llama-2-70b-chat fs-cot 0 1 1 1.000000 1.000000 False
meta-llama-3-70b-instruct baseline 3 9 12 1.732051 0.231876 False
meta-llama-3-70b-instruct zs-cot 6 19 25 2.600000 0.022882 True
meta-llama-3-70b-instruct os 4 20 24 3.265986 0.004696 True
meta-llama-3-70b-instruct os-cot 7 23 30 2.921187 0.009415 True
meta-llama-3-70b-instruct fs 0 31 31 5.567764 0.000000 True
meta-llama-3-70b-instruct fs-cot 6 6 12 0.000000 1.000000 False
meta-llama-3-8b-instruct baseline 7 32 39 4.003204 0.000337 True
meta-llama-3-8b-instruct zs-cot 10 24 34 2.400980 0.038026 True
meta-llama-3-8b-instruct os 5 35 40 4.743416 0.000019 True
meta-llama-3-8b-instruct os-cot 4 24 28 3.779645 0.000707 True
meta-llama-3-8b-instruct fs 8 29 37 3.452379 0.002308 True
meta-llama-3-8b-instruct fs-cot 4 31 35 4.563833 0.000034 True
claude-3-opus-20240229 baseline 12 26 38 2.271100 0.049984 True
claude-3-opus-20240229 zs-cot 17 22 39 0.800641 0.586163 False
claude-3-opus-20240229 os 10 16 26 1.176697 0.358975 False
claude-3-opus-20240229 os-cot 11 11 22 0.000000 1.000000 False
claude-3-opus-20240229 fs 6 9 15 0.774597 0.728687 False
claude-3-opus-20240229 fs-cot 2 4 6 0.816497 0.773438 False
claude-3-sonnet-20240229 baseline 3 23 26 3.922323 0.000431 True
claude-3-sonnet-20240229 zs-cot 3 19 22 3.411211 0.003300 True
claude-3-sonnet-20240229 os 8 18 26 1.961161 0.092843 False
claude-3-sonnet-20240229 os-cot 6 17 23 2.293659 0.072048 False
claude-3-sonnet-20240229 fs 10 17 27 1.347151 0.274523 False
claude-3-sonnet-20240229 fs-cot 5 17 22 2.558409 0.038026 True
mistral-large-latest baseline 3 6 9 1.000000 0.637718 False
mistral-large-latest zs-cot 5 6 11 0.301511 1.000000 False
mistral-large-latest os 4 11 15 1.807392 0.193859 False
mistral-large-latest os-cot 5 22 27 3.271652 0.003609 True
mistral-large-latest fs 2 6 8 1.414214 0.410773 False
mistral-large-latest fs-cot 2 5 7 1.133893 0.596799 False
Figure 15: Full experimental results for Hypothesis 3.5, framing gold vs. random ($n=200$). The perturbed problems replace reputable news agencies and universities to less trustworthy ones in syllogisms. Both the original and perturbed problems have classic patterns "All…, Some…, Some…" already removed to ensure a single confounder for token bias analysis. We run all different prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs might be less likely to be misled by less reputable names. However, such performance shifts should not happen to genuine reasoners, because the names of these entities do not affect the logical essence.
Figure 16: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.5.
D.6 Hypothesis 6
The full experimental results for Hypothesis 3.6 are shown in Figure 17, 18 and Table LABEL:table:h6.
Table 8: Full experimental results for Hypothesis 3.6
model prompting method $n_{12}$ $n_{21}$ $n^{*}$ z-stat p-value reject
gpt-3.5-turbo weak-control-zs-cot 64 423 487 16.267843 0.000000 True
gpt-3.5-turbo control-zs-cot 57 417 474 16.535348 0.000000 True
gpt-3.5-turbo weak-control-os-cot 60 250 310 10.791275 0.000000 True
gpt-3.5-turbo control-os-cot 64 230 294 9.681317 0.000000 True
gpt-4-turbo weak-control-zs-cot 8 386 394 19.043365 0.000000 True
gpt-4-turbo control-zs-cot 4 420 424 20.202746 0.000000 True
gpt-4-turbo weak-control-os-cot 6 113 119 9.808674 0.000000 True
gpt-4-turbo control-os-cot 4 126 130 10.700108 0.000000 True
gpt-4o weak-control-zs-cot 8 262 270 15.457948 0.000000 True
gpt-4o control-zs-cot 11 301 312 16.418017 0.000000 True
gpt-4o weak-control-os-cot 13 97 110 8.009086 0.000000 True
gpt-4o control-os-cot 12 112 124 8.980265 0.000000 True
llama-2-70b-chat weak-control-zs-cot 72 177 249 6.654105 0.000000 True
llama-2-70b-chat control-zs-cot 32 531 563 21.030343 0.000000 True
llama-2-70b-chat weak-control-os-cot 69 313 382 12.484126 0.000000 True
llama-2-70b-chat control-os-cot 73 429 502 15.889058 0.000000 True
meta-llama-3-70b-instruct weak-control-zs-cot 17 518 535 21.660119 0.000000 True
meta-llama-3-70b-instruct control-zs-cot 10 579 589 23.445237 0.000000 True
meta-llama-3-70b-instruct weak-control-os-cot 54 150 204 6.721344 0.000000 True
meta-llama-3-70b-instruct control-os-cot 44 174 218 8.804711 0.000000 True
meta-llama-3-8b-instruct weak-control-zs-cot 57 405 462 16.190425 0.000000 True
meta-llama-3-8b-instruct control-zs-cot 4 487 491 21.797485 0.000000 True
meta-llama-3-8b-instruct weak-control-os-cot 53 263 316 11.813423 0.000000 True
meta-llama-3-8b-instruct control-os-cot 18 301 319 15.844958 0.000000 True
claude-3-opus-20240229 weak-control-zs-cot 15 412 427 19.212177 0.000000 True
claude-3-opus-20240229 control-zs-cot 9 467 476 20.992396 0.000000 True
claude-3-opus-20240229 weak-control-os-cot 26 299 325 15.143315 0.000000 True
claude-3-opus-20240229 control-os-cot 30 212 242 11.699403 0.000000 True
claude-3-sonnet-20240229 weak-control-zs-cot 5 470 475 21.335663 0.000000 True
claude-3-sonnet-20240229 control-zs-cot 10 466 476 20.900726 0.000000 True
claude-3-sonnet-20240229 weak-control-os-cot 17 238 255 13.839557 0.000000 True
claude-3-sonnet-20240229 control-os-cot 16 250 266 14.347461 0.000000 True
mistral-large-latest weak-control-zs-cot 0 533 533 23.086793 0.000000 True
mistral-large-latest control-zs-cot 15 530 545 22.060176 0.000000 True
mistral-large-latest weak-control-os-cot 3 179 182 13.045988 0.000000 True
mistral-large-latest control-os-cot 1 209 210 14.353364 0.000000 True
Figure 17: Full experimental results for Hypothesis 3.6 ($n=800$). The perturbed problems leak hint tokens, either weak or strong hints in problem statements. We run zero-shot and one-shot prompt methods. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}<{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}.$ We conclude that LLMs still heavily rely on hint tokens for solving logical fallacy problems well.
Figure 18: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.6.